# Does Quantum Mechanics apply to Macroscopic Objects? How to define ”Macro”?

### v2

@@ Abstract
According to Quantum Mechanics a narrow wave-packet of the center of mass of any macroscopic object should smear out after some time. The problem is usually waved out by assuming that due to their heavy masses this occurs over astronomical times. Without a clear definition of macroscopic objects this remains ambiguous. On the other hand, Quantum Mechanics allows and energetically even prefers largely smeared out c.m. states that have never been seen. Why is this the real state of the world we know? Does it suggest a micro-macro threshold with different theoretical descriptions? Does micro-miniaturization offer hope for answering these questions?
@@
ORCID (https://www.orcid.org/$0$-0002-0698-4202)
@@ 1 Introduction
One should bear in mind that Quantum Mechanics is just a theoretical frame to describe the properties of a system that may be treated as being isolated and having a finite number of degrees of freedom. Its basic elements are the concept of the wave function (quantum state), the Schrödinger equation for its evolution, the observables and the probabilistic interpretation of the measurement. The system itself must be defined first from the experimental point of view and this is based on our general vague conception and long experience about the surrounding world. Then one associates a certain Hamiltonian to it and verifiable mathematical predictions of Quantum Mechanics emerge. Strictly speaking, there are no isolated physical systems and the number of degrees of freedom of a system depends on our modelling it.
Quantum Mechanics says nothing about its object or the structure of matter. It is an exceptionally successful theory, nevertheless the relationship between the macroscopic classical and the microscopic quantum mechanical description of matter is still subject of worry.
@@
Now we witness another exceptional development in producing ever smaller pieces of solid matter (semiconductor chips, micro-structures). We are approaching the microscopic world from above. Therefore, the answer to the question about the upper end of the ”quantum-domain”, might be very close.
I want here to underline some simple but often ignored aspects in this context. The theory of condensed matter is conceived as the quantum mechanical description of bound states of an enormous number of Coulomb interacting electrons and ions. This picture is very successful. In the absence of an external potential the center of mass (c.m.) motion is fully separable from the relative (internal) one [1] and its wave function satisfies the free Schrödinger equation of a particle with the total mass $M$ of the system. We apply Quantum Mechanics however, only to the relative motion, while the center of mass motion is considered irrelevant and left to the classical description.
@@
In Section 2 we recapitulate the standard quantum mechanical description of an ideal wave packet (here the free motion of the c.m), while in Section 3 we comment on the not understood absence of broad wave packets of macroscopic objects in the real word.
@@
$\psi(x,0)=\frac{1}{{(2\pi)}^{\frac{1}{4}}\sqrt{d}}e^{-\frac{x^{2}}{4d^{2}}}\qquad.$
@@
$\psi(x,t)=\frac{1}{(2\pi)^{\frac{1}{4}}\sqrt{d+\frac{\imath\hbar t}{2dM}}}e^{-\frac{x^{2}}{4(d^{2}+\frac{\imath\hbar t}{2M})}}\qquad.$
@@
$\langle x^{2}\rangle_{t}=d^{2}+(\frac{\hbar t}{2Md})^{2}\qquad.$
@@
[…]
No doubt, the c.m. of a hydrogen atom obeys Quantum Mechanics, but does it obey also the c.m of this small crystal? The answer seems to be negative. Then how do we define a ”macroscopic” object?
@@ 3 What about the ”initial” state?
For non-ideal wave-packets $\Delta x\Delta p>\hbar$ the average kinetic energy is even greater. Therefore the more smeared is the wave packet, the smaller its conserved average energy is. Although energetically preferred, no smeared out macroscopic objects have ever been seen. It looks like a ”super-selection” rule for macroscopic objects. ”God allowed them only in this state of precise c.m.?” On the other hand, a world with smeared out mass centers is hard to imagine. It won’t function. The macroscopic world, as we know it, is the only possible world!
There is another striking old problem about the application of quantum-mechanical concepts to macroscopic objects. There are no known superpositions of states of macroscopic systems. It was illustrated by the famous ”Schrödinger’s cat”.
One arrives at the conclusion that we have two different worlds: microscopic and macroscopic ones , with different theoretical descriptions.
In the above sense, I would like to encourage experiments about the imprecision of the c.m of ultra small pieces of crystals. On our current way toward micro-miniaturization we might get an answer to the question about the threshold between micro and macro. It is very plausible that the macroscopic physics does not emerge from the microscopic theory. (In a mathematical sense it is surely not the $\hbar\to 0$ limit.)
This should solve also the ”bootstrap aspect” of Quantum Mechanics i.e. the existence of macroscopic laws as prerequisite for its formulation and application. In other words: Quantum Mechanics has its limited field of validity as any other known physical theory.
@@ 4 Comments
I am aware that just touching this point might look like a heresy. In our human-centred world-image motivated by the exceptional scientific progresses of the last centuries we cheer to construct a unified picture of the world without loopholes and crevices. One looks even for ”God’s equations” forgetting that mathematics has to be connected to experiment by some interpretation and this is the most difficult one.
@@
However, all theories of physics of the past were just fragments of knowledge (models) with loose connections between them. Thermodynamics was built up much before statistical mechanics and their link is still rather vague. In the modern functional integral formalism of quantum field theory of elementary particles there is no place for wave functions, Hamiltonians and observables while its mathematical structure is highly questionable. Nevertheless, it succeeded to predict the fauna of elementary particles. However, its connection to the ordinary quantum mechanics of systems with finite degrees of freedom remains a question of belief (see also [3]). The unification of gravitation and quantum theory looks merely a dream.
@@
Theoretical physics never had a strict vertical structure. It was never derived from the ”underlying” theories of matter. I am afraid, that we are misled by our need for harmony and exaggerated confidence in our brains.

### v3

@@ Abstract
According to Quantum Mechanics a narrow wave-packet of the center of mass of any macroscopic object should spread out after some time. The problem is usually waved out by assuming that due to their heavy masses this occurs over astronomical times. Without a clear definition of macroscopic objects this remains ambiguous. On the other hand, Quantum Mechanics allows and energetically even prefers largely spread out c.m. states that have never been seen. Why is this the real state of the world we know? Does it suggest a micro-macro threshold with different theoretical descriptions? Does micro-miniaturization offer hope for answering these questions?
@@
ORCID (https://www.orcid.org/0000-0002-0698-4202)
@@ 1 Introduction
One should bear in mind that Quantum Mechanics is just a theoretical frame to describe the properties of a system that may be treated as being isolated and having a finite number of degrees of freedom. (Strictly speaking, there are no isolated physical systems and the number of degrees of freedom of a system depends on our modelling it.) The basic elements are the concept of the wave function (quantum state), the Schrödinger equation for its evolution, the observables and the probabilistic interpretation of the measurement. The system itself must be defined first from the experimental point of view and this is based on our general vague conception and long experience about the surrounding world. Then one associates a certain Hamiltonian to it and verifiable mathematical predictions of Quantum Mechanics emerge.
It is a general accepted knowledge that the hydrogen atom consist of a positively charged proton and a negatively charged electron. Its classical description could not explain neither the stability of the atom nor the discrete spectrum of emitted-absorbed light. Quantum Mechanics uses this model as a starting point for the definition of the Hamilton operator in the Schrödinger equation to get concordance with experiment, but this is far away from the picture surviving in the public’s mind.
Quantum Mechanics on its own says nothing about its object or the structure of matter. It is an exceptionally successful theory, nevertheless the relationship between the macroscopic classical and the microscopic quantum mechanical description of matter is still subject of worry.
@@
Now we witness another exceptional experimental development in producing ever smaller pieces of solid matter (semiconductor chips, micro-structures). We are approaching the microscopic world from above. Therefore, the answer to the question about the upper end of the ”quantum-domain”, might be very close.
I want here to underline some simple but often ignored aspects in this context. The theory of condensed matter is conceived as the quantum mechanical description of bound states of an enormous number of Coulomb interacting electrons and ions. This picture is very successful. In the absence of an external potential the center of mass (c.m.) motion is fully separable from the relative (internal) one [1] and its wave function satisfies the free Schrödinger equation of a particle with the total mass $M$ of the system. One applies Quantum Mechanics however, only to the relative motion, while the center of mass motion is considered irrelevant and left to the classical description.
@@
In Section 2 and 3 we recapitulate the standard quantum mechanical description of an ideal wave packet (here the free motion of the c.m), in Section 4 we state our conclusion on the not understood absence of broad wave packets of macroscopic objects in the real word while in Section 5 we formulate some general comments on physical theories.
@@
$\psi(x,0)=\frac{1}{{(2\pi)}^{\frac{1}{4}}\sqrt{d}}e^{-\frac{x^{2}}{4d^{2}}}\hskip 23.49976pt.$
@@
$\psi(x,t)=\frac{1}{(2\pi)^{\frac{1}{4}}\sqrt{d+\frac{\imath\hbar t}{2dM}}}e^{-\frac{x^{2}}{4(d^{2}+\frac{\imath\hbar t}{2M})}}\hskip 23.49976pt.$
@@
$\langle x^{2}\rangle_{t}=d^{2}+(\frac{\hbar t}{2Md})^{2}\hskip 23.49976pt.$
@@
Leaving aside the ideal limits, let us consider a numerical example for sake of illustration. Let us consider an $M=1kg$ object having an initial linear imprecision of its center of mass of d=0.5d=0.5Å. This will increase within $10^{6}$ years indeed only to 0.60.6Å.
However, for a small crystal of $M=1mg$ with the same initial imprecision of its c.m. this will increase already within 10 years to 3.283.28Å so that no crystal structure might be ”seen” any more. This seems unacceptable. Had we assumed a much smaller initial width we could get a significant broadening within minutes!
On the other hand, on a very long time scale no object may be considered as being isolated from its surrounding and the above quantum mechanical description is not any more valid. We are dealing always with open systems in weak contact with a reservoir (thermostat). This one may be ultimately the radiation field. Therefore its behaviour is turning dissipative over time (see for example [3]) and it may be described only as a mixed ensemble by the quantum-mechanical density matrix approaching equilibrium in which the non-diagonal matrix elements vanish (unless a phase transition occurs!). This asymptotic equilibrium (predicted by Thermodynamics) in the case of a free particle is a mixed state (canonical distribution) of plane waves extending into the whole space. This is not at all more satisfactory.
No doubt, the c.m. of a hydrogen atom obeys Quantum Mechanics, but does it obey also the c.m of this small macroscopic crystal? The answer seems to me negative. Then how do we define a ”macroscopic” object?
@@ 3 What about the ”initial” state?
For non-ideal wave-packets $\Delta x\Delta p>\hbar$ the average kinetic energy is even greater. Therefore the more spread is the wave packet, the smaller its conserved average energy is. Although energetically preferred, no spread out macroscopic objects have ever been seen. It looks like a ”super-selection” rule for macroscopic objects. ”God allowed them only in this state of precise c.m.?” On the other hand, a world with spread out mass centres is hard to imagine. It won’t function. The macroscopic world, as we know it, is the only possible world!
4 Conclusions
There is another striking old problem about the application of quantum-mechanical concepts to macroscopic objects. There are no known superpositions of states of macroscopic systems. It was illustrated by the famous ”Schrödinger’s cat”. In this respect it is important to mention that the superposition principle remains valid also for open quantum-mechanical systems if their density matrix obeys linear quantum-mechanical Master equations (see the old review by Gorini at al. [4] and a recent book [3] about their derivation and properties) .
An essential feature of a quantum object is his wave-like behaviour shown by diffraction. One has diffraction experiments with Buckyballs even greater than in the original experiment [5], but we cannot expect diffraction with ping-pong balls.
Another paradox, never mentioned before, follows if one admits the possibility of a full quantum-mechanical description of a macroscopic body as a superposition of states in the basis of the constituent particle states. (I mean including even its motion as a whole!) Such a state might be decomposed in a superposition of free many-body states. On the other hand, the S-matrix is defined just in terms of matrix elements of such states. A consequence of this would be that extremely high energy collisions could produce (like all kind of resonances) even macroscopic objects. It would be a very surprising event indeed!
In view of the discussion in the previous Section, experiments on the c.m. broadening of ultra small pieces of crystals (even in the Angström range!) are also relevant. On our current way toward micro-miniaturization we might get an answer to the question about the threshold between micro and macro. It is very plausible that the macroscopic physics does not emerge from the microscopic theory. In a mathematical sense it is surely not the $\hbar\to 0$ limit. (See Ref. [6] for the quasi-classical limit of Quantum Mechanics.)
A more limited interpretation in which macroscopic systems cannot be fully described in terms of quantum mechanics of their constituents, should eliminate also the ”bootstrap” aspect of Quantum Mechanics i.e. the existence of macroscopic laws as prerequisite for its formulation and application, an aspect worried many physicists.
In my understanding Quantum Mechanics has a limited domain of validity like any other known physical theory. The range of this ”quantum domain” has yet to be defined by experiments or new theories.
@@ 5 Comments
An interesting proposal for our sakes is a possible gravitational self-trapping of the c.m. wave packet. A non-linear modification of the Schrödinger equation would destroy simultaneously the superposition principle for macroscopic objects without affecting the Quantum Mechanics of microscopic ones. In the same time this might enable to define the range of validity of pure Quantum Mechanics depending on the mass and dimension of the system.
Károlyházi [9][10] (see also [11] and Penrose [12]) suggested a modification of Quantum Mechanics due to the gravitational deformation of the space-time by the constituent particles of the given body in the frame of the general relativity theory. An explicit formulation within the Newtonian frame was given by Diósi [13]and Bykov [14]. They tried to implement this idea by non-local, non-linear Schrödinger equations. An introduction of gravitational attraction between the constituent point-like particles of a system in a local way as by Coulomb interactions, would imply only irrelevant increase, respectively decrease of the Coulomb attraction or repulsion. Instead these authors identify the mass distribution of a particle to be given by the square of its wave-function. This reminds a s.c. Hartree-Hamiltonian. Th next decisive step is the identification of the inertial mass of the c.m. with its gravitational mass. Since the c.m. is just a geometrical point and not a massive particle, there is no justification at all for this assumption. Anyhow, the implementation of gravity should start within the many-particle Hamiltonian.
Diósi [13] proposes a non-local, non-linear many-body Hamiltonian. Since it is by construction translational invariant (although it is non-local, non-linear) the c.m. motion separates from the relative motion and still remains a free motion. If one includes also self-interacting terms, the c.m. motion is not separable any more. Therefore, Diosi [13] finally discusses only the self-interaction of a single particle of a given mass (the so called Schrödinger-Newton equation) , without identifying it with the center of mass of a greater system. Bykov [14] considers a quantum mechanical wave packet in the center of a macroscopic sphere with homogeneous mass distribution. Since the c.m. of such a sphere cannot be modified by his own gravitation, one has no classical motion to be quantized and again one may not speak about gravitational self-trapping of the center of mass. Leaving aside the above critics, these models lead at most to the existence of stationary states. Self-trapping should imply the stability in time of the ground state. In the Bykov [14] model it may be analytically proven that this fixed point is not an attractor. In the one-particle model of Diosi [13] only a numerical solution [16, 17] might shed light on its stability.
There is a recent review by Bassi et al. [15] about the problems of the standard Quantum Mechanics in its application to macroscopic bodies and the manifold of proposals for their solution. They insist mostly on the violation of the superposition principle by macroscopic objects and the understanding of the measurement process.
A relatively popular proposal to reconcile the collapse of the wave function during the measurement with the Schrödinger evolution is the multiple world solution of Everett [7]. It reminds me the very old paper of London and Bauer [8]. Everett introduces the second world concept instead London’s solipsist ”introspection of the observer”. The multiple world concept is however just a consistent formulation that does not modify the predictions of ordinary quantum mechanics. It does not solve the problem of the wave packet spreading.
The paper of Bassi et al. [15] is not just a critical review since my ideas about the distinct domains of quantum and classical theories overlap with theirs. Classical, macroscopic dynamics does not follow from an underlying quantum-mechanical theory. One expects a reconciliation only in a higher theory hidden behind. The necessary features of such a theory are thoroughly described therein. Another important statement by these authors I share is that ”decoherence” during the measurement cannot explain the lack of superposition of macroscopic objects. At least, the loss of coherence due to the contact with the environment (thermostat) that touches our problem of the wave packet, does not destroys the superposition principle. This follows from the linearity of quantum-mechanical Master equations for open systems as I remarked before.
@@
I am aware that these concepts are somewhat radical. In our human-centred world-image motivated by the exceptional scientific progresses of the last centuries we cheer to construct a unified picture of the world without loopholes and crevices. One looks even for ”God’s equations” forgetting that mathematics has to be connected to experiment by some interpretation and this is the most difficult one.
@@
However, all theories of physics of the past were just fragments of knowledge (models) with loose connections between them. The success of these theories was due to the possibility of simplifying some systems and phenomena as isolated ones. Thermodynamics was built up much before statistical mechanics and their link is still rather vague. Boltzmann’s famous equation, as well as hydrodynamics have never got a microscopic proof. In the modern functional integral formalism of quantum field theory of elementary particles there is no place for wave functions, Hamiltonians and observables (operators) while its mathematical structure is highly questionable. Nevertheless, it succeeded to predict the fauna of elementary particles. However, its connection to the ordinary quantum mechanics of systems with finite degrees of freedom remains a question of belief (see also [18] ). Efforts to develop a Quantum Mechanics of systems with an infinity of degrees of freedom (C* algebra) failed. The unification of gravitation and quantum theory looks merely a dream.
@@
Theoretical physics never had a strict vertical structure. It was never derived upwards from the ”underlying” theories of matter. I am afraid, that we are often misled by our need for harmony and exaggerated confidence in our brains.